# Between two descriptions of dark matter around a black hole: photon sphere, shadow, and lensing

### v1

@@ V Conclusion and discussion
We have compared two models of anisotropic dark matter (DM) distributions around a black hole (BH): the vacuum DM model and the Einstein cluster (EC) DM model. The primary physical distinction is the presence of radial pressure, which influences the time component of the metric. Vacuum DM acts as a “de Sitter” fluid and does not alter the shift function on the metric, whereas EC DM introduces an additional redshift factor whose behavior depends on the energy density profile. Consequently, the two models produce distinct signatures in electromagnetic observables. Our analysis shows that the photon sphere radius increases with DM mass in the EC model, while it shifts slightly inward for vacuum DM. Notably, EC DM exerts a significantly stronger effect on the BH shadow radius compared to vacuum DM. Lensing observables, including image positions and the differential time delay between primary and secondary images, also exhibit substantial differences between the two models.
We therefore conclude that the theoretical modeling of the DM halo surrounding a BH is crucial for determining its electromagnetic observables. If the actual DM distribution is best represented by an Einstein cluster, the vacuum DM approximation is insufficient for probing the geometry near the BH. Combining shadow measurements with strong-field lensing signatures offers a promising approach to distinguish between different DM models, particularly when supplemented by independent data from the dynamics of nearby objects. Future observational data may enable the exclusion or confirmation of competing DM descriptions.
Several caveats warrant attention. The adopted setup and assumptions may not fully capture a realistic cosmological settings. For example, time delay analyses in large scale gravitational lensing (e.g., HE 0435-1223 [43]) typically include cosmic expansion to infer the Hubble constant, which was not addressed in this study. Furthermore, lensing observables, especially time delays, are sensitive to perturbing matter both near the lens and along the line of sight. Future work that incorporates the cosmological constant into the Einstein field equations and accounts for perturbers would be valuable.
@@ Acknowledgements
We thank Byon Jayawiguna, Faris Darmawan, and Naufal Athaullah for the valuable discussions on the early stage of this work. HSR is supported by Hibah PUTI Q1 UI No. PKS-196/UN2.RST/HKP.05.00/2025.
@@
The mass function of the dark matter distribution
For a perfect fluid description of the energy–momentum tensor in a static and spherically symmetric spacetime, the mass function (also known as the Misner–Sharp mass) can be calculated by solving the Einstein field equation. In our DM case, since there is no DM mass contribution inside the cut-off radius $r_{c}$, the integration of the mass function begins at $r_{c}$, as shown in Eq. (8). The resulting mass functions for the modified Hernquist and Denhen 3/2 energy density profiles used in this work can be computed analytically, yielding
$m_{DM}^{(H)}(r)=$ $M_{DM}\frac{(r-r_{c})^{2}}{(a_{0}+r_{c})^{2}}$ (S.1)
@@
$\xi=$ $a_{0}(a_{0}-2r_{c})+\eta$
$\eta=$ $2r_{c}a_{0}\sqrt{\frac{r_{c}}{a_{0}+r_{c}}}$
where the superscripts $(H)$ and $(D)$ denote the Hernquist and Denhen 3/2 mass profiles, respectively. The total DM mass $M_{DM}$, also referred to as the Arnowitt–Deser–Misner (ADM) mass, is defined by the condition $M_{DM}=\lim_{r\to\infty}m(r)$, which is satisfied for both modified Hernquist and Denhen 3/2 profiles above.
@@
The quantities of interest for observation are then the angular image positions $\Theta^{(\pm)}$ and the time delays $t_{d}^{(\pm)}$ as functions of $\alpha$. All of these quantities can be obtained by solving the null geodesic equations given in Eqs. (10) and (11).
@@
The integration of the geodesic equation in Eq. (10) begins from the observer at $D_{OL}$ to the closest approach distance $r_{0}$—also known as the turning point, where $V(r_{0})=b^{-2}$—and is then continued to the source location at $D_{LS}$. The resulting quantity is $\phi_{end}$, which is written as
$\phi_{end}=\int_{D_{OL}}^{r_{0}}\left(\frac{d\phi}{dr}\right)dr+\int_{r_{0}}^{D_{LS}}\left(\frac{d\phi}{dr}\right)dr.$ (S.3)
Here, the impact parameter $b$ is used as the input, such that $\phi_{end}\to\phi_{end}(b)$. From this quantity, one obtains the relation between the source angle $\alpha$ and $\phi_{end}$,
@@
Additionally, the time delay $t_{d}$ is computed in a similar manner. The integration of Eq. (11) gives the elapsed time required for a photon emitted from the source to reach the observer. It is expressed as
@@
where the absolute values are used to avoid negative contributions, since the elapsed time in both segments must be positive. This provides $t_{end}(b)$. Because the distance $d$ depends on $\phi_{end}(b)$, we have $d\to d(b)$, and the time delay can be written as

### v4

@@ V Conclusion and discussion
We have compared two models of anisotropic dark matter (DM) distributions around a black hole (BH): the vacuum DM model and the Einstein cluster (EC) DM model. The primary physical distinction is the presence of negative radial pressure on the vacuum DM, which influences the time component of the metric. Vacuum DM acts as a “de Sitter” fluid and does not alter the shift function on the metric, whereas EC DM introduces an additional redshift factor whose behavior depends on the energy density profile. Consequently, the two models produce distinct signatures in electromagnetic observables. Our analysis shows that the photon sphere radius increases with DM mass in the EC model, while it shifts slightly inward for vacuum DM. Notably, EC DM exerts a significantly stronger effect on the BH shadow radius compared to vacuum DM. Lensing observables, including image positions and the differential time delay between primary and secondary images, also exhibit substantial differences between the two models.
We therefore conclude that the theoretical modeling of the DM halo surrounding a BH is crucial for determining its electromagnetic observables. If the actual DM distribution is best represented by an Einstein cluster, the vacuum DM approximation is insufficient for probing the geometry near the BH. Combining shadow measurements with strong-field lensing signatures offers a promising approach to distinguish between different DM models, particularly when supplemented by independent data from the dynamics of nearby objects—especially stars orbiting the BH, such as the S-stars around the Sagittarius A* BH [77]. Future observational data may enable the exclusion or confirmation of competing DM descriptions.
Several caveats warrant attention. The adopted setup and assumptions may not fully capture a realistic cosmological settings. For example, time delay analyses in large scale gravitational lensing (e.g., HE 0435-1223 [78]) typically include cosmic expansion to infer the Hubble constant, which was not addressed in this study. Furthermore, lensing observables, especially time delays, are sensitive to perturbing matter both near the lens and along the line of sight [79]. Future work that incorporates the cosmological constant into the Einstein field equations and accounts for perturbers from nearby galaxies and the environment along the light paths would be valuable.
Other prospective directions may also be considered. The distinct effective potentials in the EC DM and vacuum DM cases indicates that they may lead to differences in their gravitational-wave (GW) signatures, e.g., in their quasinormal modes and GW lensing echoes [10, 11, 12, 34] (see, e.g., Refs. [80, 81] for relevant discussions of GW lensing echoes). In addition, the appearance of the accretion disk around the BH may also be relevant, as we expect that EC DM and vacuum DM would yield distinct innermost circular orbits and different redshift effects on the disk intensity (see, e.g., Ref. [17]). Furthermore, one could perform a similar analysis to that presented here to compare EC DM and vacuum DM observables in their rotating configurations (e.g., those obtained numerically [65] and those derived via the Newman–Janis algorithm [9, 29, 32, 30, 31]), since realistic BHs are expected to be non-static.
@@ Acknowledgements
We thank Byon Jayawiguna, Faris Darmawan, and Naufal Athaullah for the valuable discussions on the early stage of this work. We also thank the anonymous referee for the valuable suggestions on this manuscript. HSR is supported by Hibah PUTI Q1 UI No. PKS-196/UN2.RST/HKP.05.00/2025.
@@
Deriving the metric components
For a perfect fluid description of the energy–momentum tensor in a static and spherically symmetric spacetime, the metric components can be obtained by solving the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$. In our case, the source term $T_{\mu\nu}$ is provided by the DM. We focus on the two relevant equations: $G_{0}^{0}=8\pi T_{0}^{0}$ and $G_{1}^{1}=8\pi T_{1}^{1}$. Using the line element given in Eq. (2), we obtain
$\frac{f^{\prime}}{r}+\frac{f}{r^{2}}-\frac{1}{r^{2}}$ $=-8\pi\epsilon,$ $(G_{0}^{0}=8\pi T_{0}^{0})$ (S.1)
$\frac{f\Phi^{\prime}}{r}+\frac{f^{\prime}}{r}+\frac{f}{r^{2}}-\frac{1}{r^{2}}$ $=8\pi p,$ $(G_{1}^{1}=8\pi T_{1}^{1})$ (S.2)
where denotes differentiation with respect to $r$. Equation (S.1) can be straightforwardly integrated, yielding
$f=1-\frac{2C}{r}-\frac{2m_{DM}(r)}{r},\qquad m_{DM}(r)=\int^{r}_{r_{c}}4\pi\tilde{r}^{2}\epsilon(\tilde{r})d\tilde{r}$, (S.3)
where $C$ is an integration constant and $m_{DM}(r)$ is the Misner–Sharp mass of the DM. One readily concludes that $C=M_{BH}$ by requiring a Schwarzschild geometry in vacuum ($\epsilon\to 0$). In our DM setup, since there is no DM mass contribution inside the cut-off radius $r_{c}$, the integration of the mass function begins at $r_{c}$. The function $f(r)$ then reads
$f(r)=1-\frac{2m(r)}{r},\qquad m(r)=M_{BH}+m_{DM}(r)$ (S.4)
with $m(r)$ is the total Misner-Sharp mass of the system.
The resulting mass functions for the modified Hernquist and Dehnen-3/2 energy density profiles used in this work can be computed analytically. Integrating Eq. (S.3) for both modified density profiles yields
$m_{DM}^{(H)}(r)=$ $\frac{2\pi a_{0}^{4}\epsilon_{0}}{(a_{0}+r_{c})}\frac{(r-r_{c})^{2}}{(r+a_{0})^{2}}$ (S.5)
$m_{DM}^{(D)}(r)=$ $\frac{8}{3}\pi a_{0}\epsilon_{0}\left[\frac{a_{0}\sqrt{r}[a_{0}(r-3r_{c})-2r_{c}r]}{(a_{0}+r)^{3/2}}+\eta\right]$ (S.6)
$\eta=$ $\frac{2a_{0}r_{c}^{3/2}}{\sqrt{a_{0}+r_{c}}}$
where the superscripts $(H)$ and $(D)$ denote the Hernquist and Dehnen-3/2 mass profiles, respectively. Taking the limit $r\to\infty$ of the mass function yields the Arnowitt–Deser–Misner (ADM) mass of the DM ($M_{DM}$),
$\lim_{r\to\infty}m_{DM}^{(H)}(r)=\frac{2\pi a_{0}^{4}\epsilon_{0}}{a_{0}+r_{c}}\equiv M_{DM}^{(H)},$ (S.7)
$\lim_{r\to\infty}m_{DM}^{(D)}(r)=\frac{8}{3}\pi a_{0}\xi\epsilon_{0}\equiv M_{DM}^{(H)},$ (S.8)
$\xi=a_{0}(a_{0}-2r_{c})+\eta.$
Therefore, expressing the result in terms of the total DM mass and dropping the superscripts, the mass function reads
$m_{DM}^{(H)}(r)=$ $M_{DM}\frac{(r-r_{c})^{2}}{(r+a_{0})^{2}}$ (S.9)
@@
We are now left with the $g_{tt}$ component of the metric, specifically the shift function $\Phi$. Following Ref. [68], we adopt the following ansatz for the equation of state:
$p(r)=-\epsilon(r)[1-F(r)]$ (S.12)
where $F$ is an arbitrary function. This ansatz is particularly useful for studying deviations from the (de Sitter) vacuum equation of state, for which $F=0$. In our case, we simply take $F=1$ for the EC DM and $F=0$ for the vacuum DM, so that $F(r)\to F$ is constant. Using the ansatz (S.12), substituting Eq. (S.1) into Eq. (S.2), and rearranging, we obtain
$\frac{f\Phi^{\prime}}{r}=8\pi\epsilon F.$ (S.13)
Using $f$ from Eq. (S.4), we finally obtain
$\frac{d\Phi}{dr}=\frac{8\pi r^{2}\epsilon(r)F}{r-2m(r)}.$ (S.14)
Therefore, for EC DM with $F=1$ and vacuum DM with $F=0$, we obtain Eq. (9), which describes the behavior of the shift function in both the EC DM and vacuum DM cases.
Deriving the geodesic equation for photon
The derivation of the geodesic equations in Eqs. (S.19) and (S.20) is straightforward. We begin with the general geodesic equation
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma_{\alpha\beta}^{\mu}\frac{dx^{\alpha}}{d\tau}\frac{dx^{\beta}}{d\tau}=0,$ (S.15)
with $x^{\mu}=(t,r,\theta,\phi)$ and $\tau$ an affine parameter. Restricting the motion to the equatorial plane ($\theta=\pi/2$, $\dot{\theta}=0$), we obtain the following constants of motion:
$g_{tt}\dot{t}=E,\qquad r^{2}\dot{\phi}=L,$ (S.16)
where the dotted symbols denote derivatives with respect to the affine parameter $\tau$. The four-velocity condition for null particles reads
$g_{tt}\dot{t}^{2}+g_{rr}\dot{r}^{2}+r^{2}\dot{\theta}^{2}+r^{2}\sin\theta\dot{\phi}^{2}=0.$ (S.17)
Rearranging the expression and substituting Eq. (S.16) yields
$g_{tt}g_{rr}\left(\frac{dr}{d\tau}\right)^{2}+V(r)L^{2}=E^{2},\qquad V(r)=\frac{-g_{tt}}{r^{2}}=\left[1-\frac{2m(r)}{r}\right]\frac{e^{\Phi}}{r^{2}},$ (S.18)
where $V(r)$ is the photon effective potential. Applying the chain rule to Eq. (S.18), we finally obtain
$\frac{d\phi}{dr}$ $=\frac{d\phi}{d\tau}\left(\frac{dr}{d\tau}\right)^{-1}=\pm\frac{1}{r^{2}}\left[\frac{-g_{tt}g_{rr}}{b^{-2}-V(r)}\right]^{1/2},$ (S.19)
$\frac{dt}{dr}$ $=\frac{dt}{d\tau}\left(\frac{dr}{d\tau}\right)^{-1}=\pm\frac{1}{b}\left[\frac{-g_{tt}g_{rr}}{b^{-2}-V(r)}\right]^{1/2},$ (S.20)
with $b\equiv L/E$ is known as the impact parameter. These two equations are then used to compute the photon trajectories in the spacetime, particularly for determining the lensing observables.
@@
The quantities of interest for observation are then the angular image positions $\Theta^{(\pm)}$ and the time delays $t_{d}^{(\pm)}$ as functions of $\alpha$. All of these quantities can be obtained by solving the null geodesic equations given in Eqs. (S.19) and (S.20).
@@
The integration of the geodesic equation in Eq. (S.19) begins from the observer at $D_{OL}$ to the closest approach distance $r_{0}$—also known as the turning point, where $V(r_{0})=b^{-2}$—and is then continued to the source location at $D_{LS}$. The integration yields an angle $\phi_{end}$ (see Fig. S.1),
$\phi_{end}=-\int_{D_{OL}}^{r_{0}}\left(\frac{d\phi}{dr}\right)dr+\int_{r_{0}}^{D_{LS}}\left(\frac{d\phi}{dr}\right)dr.$ (S.21)
Fixing the values of $M_{BH}$, $M_{DM}$, and $a_{0}$ gives $d\phi/dr\to d\phi(r,b)/dr$ and $r_{0}\to r_{0}(b)$. Consequently, once $D_{OL}$ and $D_{LS}$ are fixed, the integral and $\phi_{\rm end}$ depend solely on the impact parameter, so that $\phi_{\rm end}\to\phi_{\rm end}(b)$. From this quantity, one obtains the relation between the source angle $\alpha$ and $\phi_{end}$,
@@
Additionally, the time delay $t_{d}$ is computed in a similar manner. The integration of Eq. (S.20) gives the elapsed time required for a photon emitted from the source to reach the observer. It is expressed as
@@
where the absolute values are used to avoid negative contributions, since the elapsed time in both segments must be positive. Similar to Eq. (S.24), this gives $t_{\rm end}(b)$. The time required for a photon to travel a distance $d$ in flat spacetime is simply $t_{f}=cd$; since we set $c=1$, this reduces to $t_{f}=d$. The time delay is then defined as $t_{d}=t_{\rm end}-t_{f}$. Because the distance $d$ depends on $\phi_{end}(b)$, we have $d\to d(b)$, and the time delay can be written as
@@
Multiplication of secondary images
To illustrate the origin of multiple secondary images, we perform a backward ray-tracing procedure, computing photon trajectories from $D_{OL}/M_{BH}=10^{11}$ toward the BH and extending to infinity for a range of representative impact parameters. In the physical interpretation, these light rays are emitted from arbitrary sources toward the BH and then propagate to the observer’s location. For this setup, we use the Hernquist EC DM with $M_{DM}/M_{BH}=100$ and $a_{0}/M_{BH}=10^{4}$. The resulting photon trajectories are shown in Fig. S.2.
Figure S.2: Photon trajectories originating from arbitrary sources on the left side of the figure, propagate toward the BH, and then continue to the observer at $(X/M_{BH},Y/M_{BH})=(10^{11},0)$ with impact parameters $b/M_{BH}\in[4\times 10^{2},4\times 10^{4}]$. The BH is located at $(X,Y)=(0,0)$, indicated by the small solid dot. The size of the solid dot does not represent the physical size of the BH. Darker trajectories correspond to smaller impact parameters (i.e., photons passing closer to the BH), while brighter trajectories have larger impact parameters. The thicker curves highlight representative photon trajectories that lead to image multiplication: light emitted from the yellow star is received by the observer along two distinct photon paths with different impact parameters, and therefore appears at two different observation angles.
One might notice the sharp turn of light rays occurring at $X/M_{BH}\sim 0$. This happens because the plot scale is extremely large and highly unequal, being of order $\sim 10^{6}M_{BH}$ on the horizontal ($x$-)axis and $\sim 10^{4}M_{BH}$ on the vertical ($y$-)axis. As a result, small changes along the $y$-axis appear significantly larger than those along the $x$-axis. The clear, smooth bending of light, as illustrated in Fig. S.1, occurs much closer to the BH horizon at $(X,Y)=(0,0)$, at radii satisfying $\sqrt{X^{2}+Y^{2}}=2M_{BH}$. Therefore, at the scale of the plot, the light bending appears as a sharp turn at $X/M_{BH}\sim 0$. This choice of scale is made solely to illustrate more clearly the origin of image multiplication of a source, e.g., the yellow star.
It is evident that there exist specific ranges of smaller impact parameters (darker red curves) whose deflection is weaker than that of some photons with larger impact parameters (brighter orange–yellow curves). Consequently, photons belonging to these lower-impact-parameter ranges can intersect those with higher impact parameters at certain locations away from the BH (e.g., at the position of the yellow star in Fig. S.2). Since the impact parameter is directly related to the observation angle $\Theta^{(\pm)}$ (cf. Eq. (S.23)), the same source can be detected at (at least) two distinct angles. This leads to the formation of multiple (secondary) images. We did not investigate this behavior in further detail, as it lies beyond the scope of the present study.